OpenStudy (anonymous):

HELP! I don't know how to solve -3x + 4 <= x + 8 can someone please explain how to solve it that would be so helpful

OpenStudy (turingtest):

\[-3x+4 \ge x+8\]just pretend it's an equals sign, and combine like terms (try to keep x positive if possible):\[4-8 \ge x+3x\]\[-4 \ge 4x\]\[-1 \ge x\]

OpenStudy (anonymous):

thank you!